show the calculation for the mass of CaCL2 in the unknown mixture

please answer both questions.
Answer #1

Mass of CaCl2 taken = 108.1 - 107.5 = 0.6g

Number of moles of CaCl2 = 0.6g/110.98g/mol = 0.00540638mol

Number of moles of Calcium in Ca3(PO4)2 = number of moles of CaCl2 ÷ 3 = 0.00540638mol /3 =0.00180213mol

Mass of Ca3(PO4)2 ppt formed = 0.00180213mol × 310.177g/mol = 0.55898g (theoretical yield)

% yield =( experimental yield/ theoretical yield )×100 %

= 0.6g/0.55898g × 100% = 107.3 %

= 107.3%

( The yield is greater than 100% means the ppt might not be dried properly . Or there are some error in measuring of mass of ppt).